On a recent quiz, the class mean was 73 with a standard deviation of 4. Calculate the z-score (to 4 decimal places) for a person who received score of 87.

z-score:

Is this unusual?

  • Unusual
  • Not Unusual
Answer #1

Z score = (x-mu)/sigma =(87 - 73)/4 = 3.5

unusual.

for further any queries please comment and thank you.